A 112 kg man stands 551 m from a building of 25,500 kg. What is the gravitational force of the man due to the building? What is the gravitational force of the building due to the man?
1
Expert's answer
2020-12-10T11:02:31-0500

"F = \frac {-G*m*M}{r^2}"

Both the bodies with masses m and M exert same force F on each other.

The negative sign implies Attractive Nature of the force.

F ="- 6.67 *\frac{ 112*25500}{551^2}" *10-11 N

F = -6.275 * 10-11 N (approx)